OpenStudy (anonymous):

Situation: A farmer in China discovers a mammal hide that contains 54% of it's original amount of c-14. Find the age of the mammal hide to the nearest year.

OpenStudy (anonymous):

this sounds like biology or something.

OpenStudy (anonymous):

it's math.

OpenStudy (anonymous):

Ok well maybe I just don't get it... What class is it for?

OpenStudy (anonymous):

algebra 2

OpenStudy (anonymous):

really? is there any other information with the problem?

OpenStudy (anonymous):

\[N = N _{0}e ^{-kt}\]

OpenStudy (anonymous):

N is amout remaining now; and No is initial amount?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so, N =54, No =100 and what is k?

OpenStudy (anonymous):

54/100 = e^(-k *t) log(.54) =-k *t if u were given the value of k, just plug in and findout t.

OpenStudy (anonymous):

k = .0001

OpenStudy (anonymous):

plug in log(.54) = -0.0001 * t log(.54) / -0.0001 = t

OpenStudy (anonymous):

i dont hav the calculator with me now, u have to feed it in and see what T is.

OpenStudy (anonymous):

log (.54) is a -ve value / - value of that k = +ve time...so thats right